BluStar Company has two service departments, Administration and Accounting, and two operating departments, Domestic and International. Administration costs are allocated on the basis of employees, and Accounting costs are allocated on the basis of number of transactions. A summary of BluStar operations follows.

Administration Accounting Domestic International
Employees 28 46 26
Transactions 40,000 24,000 96,000
Department direct costs $ 353,000 $ 146,000 $ 945,000 $ 3,640,000

Required:

a. Allocate the cost of the service departments to the operating departments using the direct method.
b. Allocate the cost of the service departments to the operating departments using the step method. Start with Administration.
c. Allocate the cost of the service departments to the operating departments using the reciprocal method.

Required A:

Allocate the cost of the service departments to the operating departments using the direct method. (Do not round intermediate calculations. Negative amounts should be indicated by a minus sign.)

To
From Administration Accounting Domestic International
Department costs $353,000 $146,000 $945,000 $3,640,000
Administration allocation (353,000)
Accounting allocation (146,000)
Total cost $0 $0 $945,000 $3,640,000

Required B:

Allocate the cost of the service departments to the operating departments using the step method. Start with Administration. (Do not round intermediate calculations. Negative amounts should be indicated by a minus sign.)

To
From Administration Accounting Domestic International
Department costs
Administration allocation
Accounting allocation
Total cost $0 $0 $0 $0

Required C:

Allocate the cost of the service departments to the operating departments using the reciprocal method. (Do not round intermediate calculations. Negative amounts should be indicated by a minus sign. Round your final answers to the nearest whole dollar amounts.)

To
From Administration Accounting Domestic International
Department costs
Administration allocation
Accounting allocation
Total cost $0 $0 $0 $0
Answer #1
B. To
From Administration Accounting Domestic International
Department costs 353000 146000 945000 3640000
Administration allocation -353000 98840 162380 91780
(On the basis of employees at 28:46:26 or 14:23:13) 353000*(14/50) 353000*(23/50) 353000*(13/50)
Accounting allocation -244840 48968 195872
(On the basis of transactions at 24000:96000 or 1:4) 244840*(1/5) 244840*(4/5)
Total cost 0 0 1156348 3927652
c. Under reciprocal method it is assumed that cost from other service departments is included in the cost of individual service departments
Total cost of a service department=Direct cost of the service department+Cost allocated to the service department
Let's assume that cost of administration=D1
Cost allocated from the accounting on the basis of transactions
Transactions for administration=40000
Total transaction=40000+24000+96000=160000
Ratio for allocation to administration=40000/160000=0.25
D1=353000+0.25 D2 …………. 1
Let's assume that cost of accounting=D2
Cost allocated from the administration on the basis of employees
Employees for accounting=28
Total employees=28+46+26=100
Ratio for allocation to accounting=28/100=0.28
D2=146000+0.28 D1 …………. 2
Apply 1 in 2
D2=146000+0.28*(353000+0.25D2)
D2=146000+98840+0.07D2
D2=244840+0.07D2
D2-0.07D2=244840
0.93D2=244840
D2=244840/0.93=$ 251881.7=$ 263268.8=$ 263269
Apply D2 in 1
D1=353000+(0.25*263269)=353000+65817.25=$ 418817.3=$ 418817
Allocation:
From Administration Accounting Domestic International
Department costs 353000 146000 945000 3640000
Administration allocated -418817 117269 192656 108892
(On the basis of employees at 28:46:26 or 14:23:13) (418817*14/50) (418817*23/50) (418817*13/50)
Accounting allocation 65817 -263269 39490 157961
(On the basis of transactions at 40000:24000:96000 or 5:3:12) (263269*5/20) (263269*3/20) (263269*12/20)
Total cost allocated 0 0 1177146 3906854